You want to accumulate $28,000 in 10 years using annual payments, the first one in a year. If the first four payments are $1500 each, what will be the annual payment for the next six years if the effective annual return is 10% for the entire period?
Answer #1

Rate =10%
Annual Payment for first 4 years =1500
FV of first 4 payments at end of year 4 =PMT*(((1+r)^n-1)/r) =1500*((1+10%)^4-1)/10%) =6961.50
FV of this payment at end of 10 years =6961.50*(1+10%)^6 =12332.72

The FV of remaining 6 payments =28000-12332.72 =15667.28
Number of Periods =6
The annual payment =FV/((1+r)^n-1)/r) =15667.28/((1+10%)^6-1)/10%) =2030.60